assuming that the following tables are for linear functions, fill in the blank outputs and find formulas for these functions
Input Output I guessed
0 __ 4
1 __ 5.5
2 7
3 __ 8.5
4 10
5 __ 11.5
6 __ 13
Not sure if output is right if it is not sure how to get the formula

Answer by jim_thompson5910(35256)

You have the correct outputs. As x increases by 2, then y increases by 3. So as x increases by 1, y increases by 1.5 (divide each value in half)

So the slope is m = (Change in y)/(Change in x) = 1.5/1 = 1.5

So m = 1.5

The y-intercept is (0, 4)


So the equation is y = 1.5x + 4


Note: the equation above is in the form y = mx+b where m is the slope and (0, b) is the y-intercept.


You can check this by plugging in x = 2 to get y = 7. Also you can plug in x = 4 to get y = 10